Dr. P’s grandfather clock has a pendulum of length of 1.0 cm. Determine: (a) its period on earth, (b) its
period on the moon where g = 1.62 m/s2.
1
Expert's answer
2020-10-15T03:18:47-0400

Solution

time period is given by

"T=2\pi \sqrt{\frac{l}{g}}"

a) time period on earth

"T_e=2\times3.14\sqrt{\frac{0.01}{9.8}}=0.032s"

b) time period on moon

"T_e=2\times3.14\sqrt{\frac{0.01}{1.62}}=0.078s"